OpenStudy (anonymous):

given y=csc^2 x. what is the instantaneous rate of change at x = pie/2

jimthompson5910 (jim_thompson5910):

first find dy/dx or y'

OpenStudy (anonymous):

-cscx*cotx?

jimthompson5910 (jim_thompson5910):

close, that's the derivative of just csc(x)

jimthompson5910 (jim_thompson5910):

but you want the derivative of csc^2(x)

OpenStudy (anonymous):

i don't know how to do that

jimthompson5910 (jim_thompson5910):

use the chain rule

jimthompson5910 (jim_thompson5910):

y=csc^2 x dy/dx=2*csc x*d/dx[csc(x)] dy/dx=2*csc x*(-csc(x)*cot(x)) dy/dx = -2*csc^2 x*cot(x)

jimthompson5910 (jim_thompson5910):

now plug in x = pi/2 to find the instantaneous rate of change at x = pi/2

OpenStudy (anonymous):

is that it?

jimthompson5910 (jim_thompson5910):

yes, well you need to plug in x = pi/2 to get a value, but pretty much